# Data Quality Issues in Multilingual Speech Datasets: The Need for Sociolinguistic Awareness and Proactive Language Planning

### v1

@@ Abstract
Our quality audit for three widely used public multilingual speech datasets—Mozilla Common Voice 17.0, FLEURS, and VoxPopuli—shows that in some languages, these datasets suffer from significant quality issues. We believe addressing these issues will make these datasets more useful as training and evaluation sets, and improve downstream models. We divide these quality issues into two categories: micro-level and macro-level. We find that macro-level issues are more prevalent in less institutionalized, often under-resourced languages. We provide a case analysis of Taiwanese Southern Min (nan_tw) that highlights the need for proactive language planning (e.g. orthography prescriptions, dialect boundary definition) and enhanced data quality control in the process of Automatic Speech Recognition (ASR) dataset creation. We conclude by proposing guidelines and recommendations to mitigate these issues in future dataset development, emphasizing the importance of sociolinguistic awareness in creating robust and reliable speech data resources.
@@ 5.3 Root Causes of the nan_tw Data Issues
To understand the rationale behind the dictionary-style structure and dual-script representation in MCV17 nan_tw, we contacted the voice and text contributors. They explained that most TSM speakers are not proficient in reading and writing TSM. Moreover, among those who are literate in the language, preferences and proficiency levels for different writing standards vary considerably. Some are only comfortable with Sinographs, others with Romanization, while some prefer a mixture of both. To maximize participation and facilitate data collection, the contributors opted to include both Sinographs and Romanization, so that all potential contributors could read the prompts.
@@ 7.1 Speech Dataset as a tool for Language Planning
Notably, we recommend investigating how the process of community-driven ASR dataset construction can assist language revitalization and corpus planning efforts, particularly in the orthography development of less-institutionalized languages Bianco (2018). Markl and McNulty (2022) pointed out that ASR datasets can be infrastructures for language communities, while current maintenance and curation of speech datasets for uninstitutionalized languages are still lacking. Reitmaier et al. (2024) offers a example of enabling speech technologies for unwritten languages. However, how speech dataset construction can facilitate the development of writing system still remains to be explored.

### v2

@@ Abstract
Our quality audit for three widely used public multilingual speech datasets—Mozilla Common Voice 17.0, FLEURS, and VoxPopuli—shows that in some languages, these datasets suffer from significant quality issues, which may obfuscate downstream evaluation results while creating an illusion of success. We divide these quality issues into two categories: micro-level and macro-level. We find that macro-level issues are more prevalent in less institutionalized, often under-resourced languages. We provide a case analysis of Taiwanese Southern Min (nan_tw) that highlights the need for proactive language planning (e.g. orthography prescriptions, dialect boundary definition) and enhanced data quality control in the dataset creation process. We conclude by proposing guidelines and recommendations to mitigate these issues in future dataset development, emphasizing the importance of sociolinguistic awareness and language planning principles. Furthermore, we encourage research into how this creation process itself can be leveraged as a tool for community-led language planning and revitalization.
@@ 5.3 Root Causes of the nan_tw Data Issues
To understand the rationale behind the dictionary-style structure and dual-script representation in MCV17 nan_tw, we contacted the voice and text contributors. They explained that most TSM speakers are not proficient in reading or writing TSM. Moreover, among those who are literate in the language, preferences and proficiency levels for different writing standards vary considerably. Some are only comfortable with Sinographs, others with Romanization, while some prefer a mixture of both. To maximize participation and facilitate data collection, the contributors opted to include both Sinographs and Romanization, so that all potential contributors could read the prompts.
@@ 7.1 Speech Dataset as a tool for Language Planning
The creation of crowdsourced datasets like Common Voice offers a unique microcosm of community-based language planning in action. This process inherently surfaces and forces negotiation on unresolved questions of orthography, lexicon, and dialectal norms. We argue that this phenomenon should not be viewed as an obstacle, but as a critical opportunity. Such dataset construction, particularly for less-institutionalized languages, should be re-conceptualized as not merely data collection, but a live, community-led language planning initiative. As Markl and McNulty (2022) argue, ASR datasets are vital "infrastructures" for language communities, but their maintenance and curation are often lacking.
We argue that a critical next step for the research community is to explore how to deliberately leverage this process. Learning from Reitmaier et al. (2024) who enables speech technologies even for unwritten languages, future work should focus on developing methodologies and frameworks that empower communities to use dataset creation as an explicit tool for grassroots language planning and revitalization.